A constable follow a thief who is 200m ahead of him. If the constable and the thief runs at speeds of 8km/hr and 7km/hr respectively,how soon would the constable catch the thief?

Grade:7

1 Answers

Arun
6 years ago
Let c=constable,  t =theif
Veocity:v(c)= 8km/hr=8*5/18 m/sec=40/18 m/sec
Similarly, v(t)=35/18 m/sec
As the velocity is uniform, we can use the formula,dist=v.t
Let constable overtakes the theif at a dist. d from the theif, so 
the time taken to meet at that point will be same for both of them.
i.e. t(c)=t (t)
Therefore, (200+d)/v (c)=d/v(t)         ( time= dist/v  and constable has to cover 200m more than theif)
(200 + d) * 18/40 = d *18/ 35
 by putting the values  of velocities and solving we get d=1400 m